OpenStudy (anonymous):

help pleasee!! What is the equation of the line that is perpendicular to y = 2x – 4 and has a y-intercept of –2?

OpenStudy (anonymous):

@TuringTest

OpenStudy (sheraz12345):

m=-1/2 and c=-2 Put values in y=mx+c

OpenStudy (anonymous):

LIKE -1/2x+-2

OpenStudy (anonymous):

-1/2x+2

OpenStudy (sheraz12345):

-1/2x-2

OpenStudy (anonymous):

can u help me on this one too The slope of a line is –4 and its y-intercept is (0, 4). What is the equation of the line that is perpendicular to the first line and passes through (–8, 2)?